# RegNLP in Action: Facilitating Compliance Through Automated Information Retrieval and Answer Generation

### v1

@@ Abstract
Regulatory documents, issued by governmental regulatory bodies, establish rules, guidelines, and standards that organizations must adhere to for legal compliance. These documents, characterized by their length, complexity and frequent updates, are challenging to interpret, requiring significant allocation of time and expertise on the part of organizations to ensure ongoing compliance.Regulatory Natural Language Processing (RegNLP) is a multidisciplinary subfield aimed at simplifying access to and interpretation of regulatory rules and obligations. We define an Automated Question-Passage Generation task for RegNLP, create the ObliQA dataset containing 27,869 questions derived from the Abu Dhabi Global Markets (ADGM) financial regulation document collection, design a baseline Regulatory Information Retrieval and Answer Generation system, and evaluate it with RePASs, a novel evaluation metric that tests whether generated answers accurately capture all relevant obligations and avoid contradictions.
RegNLP in Action: Facilitating Compliance Through Automated Information Retrieval and Answer Generation
@@ 1 Introduction
Regulatory documents are formal quasi-legal texts issued by governmental regulatory bodies that set out rules, guidelines, and standards that organizations must follow to ensure compliance. These documents typically cover a wide range of topics, from environmental and financial compliance to workplace safety and data protection. Characterized by their complex and precise language, they are designed to cover all potential scenarios, but can be challenging to interpret without specialized knowledge. Additionally, regulatory documents are frequently updated to reflect new laws, technological changes, or shifts in societal norms, necessitating continuous monitoring and adaptation by organizations to meet new compliance requirements. Compliance with these regulations requires considerable investment in terms of time and expertise. For example, between 1980 and 2020, the U.S. public and private sectors spent an estimated 292.1 billion hours complying with 36,702 regulations, accounting for approximately 3.2% of total annual working hours Kalmenovitz (2023). Moreover, compliance errors can lead to severe penalties, as, for example, demonstrated by the recent C\$7.5 million fine levied against the Royal Bank of Canada by FINTRAC for failing to report suspicious transactionsBloomberg (2023). Regulatory Natural Language Processing (RegNLP) is a recent multidisciplinary subfield designed to simplify access to and interpretation of regulatory rules and obligations, reducing errors and enhancing operational efficiency.
@@
    Automated Question-Passages Generation: We introduce a framework that employs Large Language Model (LLM) agents to automatically generate questions, with Natural Language Inference (NLI) integrated at the validation stage. This framework can be applied to specialized regulatory document collections to create datasets for the development and evaluation of question-answering and passage retrieval systems.
@@
    ObliQA: We have developed a multi-document, multi-passage QA dataset  [^1] https://github.com/RegNLP/ObliQADataset specifically developed to facilitate research in RegNLP. It consists of 27,869 questions and their associated source passages, all derived from the full collection of regulatory documentation provided by Abu Dhabi Global Markets (ADGM) [^2] https://www.adgm.com/, the authority overseeing financial services in the UAE’s free economic zones.
@@
    Regulatory Information Retrieval and Answer Generation: We introduce a task that first involves passage retrieval given a specific regulatory query, identifying all relevant obligations distributed across all regulatory documents. Following this, the task requires generating comprehensive answers based on the retrieved passages. These answers must synthesize information from multiple sources, ensuring that all pertinent obligations are included to fully address the regulatory query. The goal is to create responses that are both complete and concise, providing clear guidance without overwhelming the user with extraneous information.
@@
    Regulatory Passage Answer Stability Score, We develop RePASs a novel reference-free metric aimed at evaluating the quality of generated answers in regulatory compliance contexts. RePASs is tailored to the complex nature of regulatory texts, ensuring that all relevant obligations are accurately captured, fully reflected in the answers, ideally devoid of contradictions.
This paper is structured as follows. Section 2 discusses existing work on RegNLP. Section 3 describes the creation of the ObliQA dataset, and Section 4 introduces the RePASs evaluation metric. Section 5 describes our 2-step baseline Regulatory Information Retrieval and Answer Generation (RIRAG) system, which first retrieves relevant passages for a given query and secondly generates an answer from these passages, and presents our evaluation of both steps. Section 6 concludes the paper and provides directions for future work. Section Limitations addresses the limitations of this work.
@@ 2.1 Research on Regulatory NLP
These studies highlight the potential of recent NLP techniques and tools for improving the retrieval and extraction of regulatory information. Our work extends RegNLP by firstly developing a novel dataset in the financial domain from the complete document collection of a single regulatory authority, and secondly defining a novel regulatory question-answering (QA) task and evaluation metric.
2.2 Synthetic Data Generation with LLMs
Recent research studies have demonstrated the versatile capabilities of LLMs in enriching QA datasets across diverse domains. Alberti et al. (2019) pioneered methods for generating synthetic QA corpora, using roundtrip consistency for validation and achieving state-of-the-art performance on tasks such as SQuAD2 and NQ. Maatouk et al. (2023) introduced a zero-shot learning approach for neural passage retrieval, using synthetic question generation and hybrid term-neural models to enhance retrieval performance without extensive domain-specific data. Abdullin et al. (2023) focused on synthetic dialogue dataset generation, automating dialogues for linear programming problems using LLMs, thereby aiding in training conversational agents. Maatouk et al. (2023) developed TeleQnA, specifically evaluating LLMs’ telecommunications knowledge, while Tihanyi et al. (2024) introduced CyberMetric for assessing cybersecurity knowledge. These datasets enable evaluation and further training of LLMs to enhance their domain-specific capabilities. Additionally, multilingual question generation for health education, as explored by Ackerman and Balyan (2024), highlights the effectiveness of LLMs in diverse linguistic contexts. Collectively, these studies underscore the transformative potential of LLMs in refining QA datasets tailored to specific domains, fostering the development of more specialized and effective AI tools.
@@ 2.3 Retrieval-Augmented Generation
RegNLP can benefit from advances in retrieval-augmented generation (RAG) Lewis et al. (2020) and other related technologies to simplify access to regulatory documents and improve compliance. Regulatory documents, known for their complexity and frequent updates, pose unique challenges that can be mitigated through methods that leverage the strengths of NLP and information retrieval systems. RAG has significantly improved the accuracy, efficiency, and trustworthiness of LLMs by integrating external, contextually relevant and up-to-date information. Notable approaches include: Self-RAG Asai et al. (2023) enhances response quality by incorporating self-reflection mechanisms; RAFT Zhang et al. (2024) optimizes domain-specific knowledge retrieval by filtering out irrelevant data and improving reasoning chains; RobustRAG Xiang et al. (2024) and RQ-RAG Chan et al. (2024) tackle challenges such as retrieval corruption and ambiguous queries; PipeRAG Jiang et al. (2024a) improves performance by reducing retrieval latency. However, these studies do not consider regulatory documents so we are interested in testing the ability of RAG methods for solving the QA task for regulatory questions.
@@ 3 ObliQA: Obligation Based Question Answering Dataset
Figure 1: AI-driven Framework for Generating Question-Passage Pairs to Create ObliQA: Obligation-based Question-Answering Dataset for Regulatory Compliance. This flowchart presents the process from data collection and standardization to question generation and NLI-based validation for constructing a synthetic dataset.
@@ 3.1 Data Collection
The regulatory documents provided by ADGM are in .docx format, each ranging from 30 to 100 pages and written in a broadly legalistic style. These documents feature complex internal structures, including numerous (sub)subsections, numbered clauses, and cross-references within and across documents. In total, there are 40 documents comprising approximately 640,000 words.

### v2

# RIRAG: Regulatory Information Retrieval and Answer Generation

@@ Abstract
Regulatory documents, issued by governmental regulatory bodies, establish rules, guidelines, and standards that organizations must adhere to for legal compliance. These documents, characterized by their length, complexity and frequent updates, are challenging to interpret, requiring significant allocation of time and expertise on the part of organizations to ensure ongoing compliance. Regulatory Natural Language Processing (RegNLP) is a multidisciplinary field aimed at simplifying access to and interpretation of regulatory rules and obligations. We introduce a task of generating question-passages pairs, where questions are automatically created and paired with relevant regulatory passages, facilitating the development of regulatory question-answering systems. We create the ObliQA dataset, containing 27,869 questions derived from the collection of Abu Dhabi Global Markets (ADGM) financial regulation documents, design a baseline Regulatory Information Retrieval and Answer Generation (RIRAG) system and evaluate it with RePASs, a novel evaluation metric that tests whether generated answers accurately capture all relevant obligations while avoiding contradictions.
RIRAG: Regulatory Information Retrieval and Answer Generation
@@ 1 Introduction
Regulatory documents are formal semi-legal texts issued by governmental regulatory bodies, setting out rules, guidelines, and standards organizations must follow to ensure compliance. These documents typically cover a wide range of topics, from environmental and financial compliance to workplace safety and data protection. Characterized by their complex and precise language, they are designed to cover all potential scenarios, but can be challenging to interpret without specialized knowledge. Additionally, regulatory documents are frequently updated to reflect new laws, technological changes, or shifts in societal norms, necessitating continuous monitoring and adaptation by organizations to meet new compliance requirements. Compliance with these regulations requires considerable investment in terms of time and expertise. For example, between 1980 and 2020, the U.S. public and private sectors spent an estimated 292.1 billion hours complying with 36,702 regulations, accounting for approximately 3.2% of total annual working hours Kalmenovitz (2023). Moreover, compliance errors can lead to severe penalties, as, for example, demonstrated by the recent C\$7.5 million fine levied against the Royal Bank of Canada by FINTRAC for failing to report suspicious transactions Bloomberg (2023). Regulatory Natural Language Processing (RegNLP) is a recent multidisciplinary field designed to simplify access to and interpretation of regulatory rules and obligations, reducing errors and enhancing operational efficiency.
Question-answer systems for regulatory documents enable stakeholders to better navigate complex rules and obligations. Stakeholders often need to extract precise obligations from these documents to address queries like, "What steps must we take when a transaction is suspected of involving money laundering?". However, due to the intricate structure of regulatory texts and the importance of adhering to semi-legal rules, generating accurate and comprehensive answers requires careful retrieval and synthesis of relevant obligations. These answers must include all necessary obligations while avoiding omissions or contradictions (see Appendix A.1). This challenge highlights the need for advanced systems designed to automate and improve this process.
@@
    Automated Question-Passages Generation: We introduce a framework that employs Large Language Model (LLM) to automatically generate questions, with Natural Language Inference (NLI) integrated at the validation stage. This framework can be applied to specialized regulatory document collections to create datasets for the development and evaluation of question-answering and single passage or multi passages retrieval systems.
@@
    ObliQA - Obligatory Question-Answer dataset: We develop a multi-document, multi-passage QA dataset [^1] https://github.com/RegNLP/ObliQADataset specifically to facilitate research in RegNLP. It consists of 27,869 questions and their associated source passages, all derived from the full collection of regulatory documentation provided by Abu Dhabi Global Markets (ADGM) [^2] https://www.adgm.com/, the authority overseeing financial services in the Emirates of Abu Dhabi international financial centre and free zone.
@@
    RIRAG - Regulatory Information Retrieval and Answer Generation: We introduce a task that first involves passage retrieval given a specific regulatory query, identifying all relevant obligations distributed across all regulatory documents. Following this, the task requires generating comprehensive answers based on the retrieved passages. These answers must synthesize information from multiple sources, ensuring that all pertinent obligations are included to fully address the regulatory query. The goal is to create responses that are both complete and concise, providing clear guidance without overwhelming the user with extraneous information.
@@
    RePASs - Regulatory Passage Answer Stability Score: We propose RePASs [^3] https://github.com/RegNLP/RePASs, a novel reference-free metric designed to evaluate the quality of generated answers in regulatory compliance contexts. RePASs is tailored to the complex nature of regulatory texts, ensuring that all relevant obligations are accurately captured, fully reflected in the answers, and ideally devoid of contradictions. To effectively ensure that all obligations are covered, we developed an obligation classifier [^4] https://github.com/RegNLP/ObligationClassifier model fine-tuned on LegalBERT Chalkidis et al. (2020), optimized for the detection of obligatory sentences within regulatory passages.
@@ 2.1 Research on Regulatory NLP
Despite these advances, regulatory document retrieval faces challenges such as implicit obligations, multi-document queries, and cross-referenced legal clauses. Current systems often struggle to synthesize comprehensive answers when obligations are dispersed or fragmented, and they inadequately address contradictions and obligation coverage. Our work addresses these gaps by introducing a novel dataset and QA task tailored to the financial domain, built from the comprehensive collection of ADGM regulatory documents. Unlike previous datasets that focus on narrower aspects of compliance, our dataset captures a broader range of obligations across multiple documents. We also propose the RePASs evaluation metric, which goes beyond relevance to assess obligation coverage and contradictions, areas where existing metrics fall short.
@@ 2.2 Retrieval-Augmented Generation
RegNLP can benefit from advances in retrieval-augmented generation (RAG) Lewis et al. (2020) and other related technologies to simplify access to regulatory documents and improve compliance. Regulatory documents present two key challenges for application systems: (1) frequent updates that can quickly render systems outdated, and (2) the complexity of specialized terminology and cross-document references that demand a comprehensive understanding of the entire document collection. RAG has significantly improved the accuracy, efficiency, and trustworthiness of LLMs by integrating external, contextually relevant and up-to-date information. Notable approaches include: Self-RAG Asai et al. (2023) enhances response quality by incorporating self-reflection mechanisms; RAFT Zhang et al. (2024) optimizes domain-specific knowledge retrieval by filtering out irrelevant data and improving reasoning chains; RobustRAG Xiang et al. (2024) and RQ-RAG Chan et al. (2024) tackle challenges such as noisy retrieval results and ambiguous queries; PipeRAG Jiang et al. (2024a) improves performance by reducing retrieval latency. However, these studies do not consider regulatory documents so we are interested in testing the ability of RAG methods for solving the QA task for regulatory questions.
@@ 3 ObliQA: Obligation Based Question Answering Dataset
Figure 1: AI-driven framework for generating question-passages pairs to create ObliQA: Obligation-Based Question Answering dataSet for regulatory compliance. This flowchart presents the process from data collection and standardization to question generation and NLI-based validation for constructing a synthetic dataset.
@@ 3.1 Data Collection
The regulatory documents provided by ADGM consist of 40 documents totaling 640k words, each ranging from 30 to 100 pages. These documents are carefully structured with numbered clauses and subsections to maintain a hierarchical format. More details on document standardization and structuring processes can be found in the Appendix A.2.
3.1.1 Question Generation
Regulatory documents contain extensive descriptions and titles, but for question generation, only passages containing obligations are considered essential, ensuring the questions are grounded in core compliance requirements.
Input Definition:
Inputs for Multi-Passage Questions: For broader queries, inputs encompass multiple source passages. Sixty-nine distinct topics related to finance and regulatory compliance (see Appendix A.3 for a sample) guide the aggregation of related regulations into meaningful groups. These groups then form the basis for multi-passage questions by randomly sub-grouping a specified number of rules per question.
For single passage questions, the prompt is:
#Questions #Passages Relevant Irrelevant Null Relevant but Indirect
Group 1 79 79 78 1 0 0
Group 2 178 280 194 37 2 47
Group 3 230 428 356 14 6 52
Group 4 245 490 471 7 0 12
Group 10 259 570 506 6 0 58
Total 991 1847 1605 65 8 169
Table 2: Expert evaluation of passage relevance in the ObliQA dataset, categorized by groups.
3.2 Question-Passage(s) Validation via NLI
Validating the generated question-passage(s) pairs using Natural Language Inference (NLI) involves assessing the semantic relationship between each question and its corresponding source passage(s) to ensure alignment Dagan et al. (2006). The validation process uses the nli-deberta-v3-xsmall model with passage(s) as the premise and question as the hypothesis. For each passage associated with a question, NLI is applied to evaluate the relationship. The validation results are categorized as follows.
  • Entailment: The passage entail the questions being retained.
  • Contradictions: The passage that contradict the question is eliminated.
  • Neutral: The neutral passage is retained if the absolute difference between its neutral score and the entailment score is smaller than the difference between its neutral score and the contradiction score; otherwise, it is discarded.
The final dataset consists of approximately 21k questions linked to single passages and 5k questions linked to multiple passages. The distribution of questions among different groups, indicating the number of input passages, is detailed Table 1. The further distribution between the training, testing and development sets is shown in Appendix A.4. A sample of the dataset is provided in Appendix A.5.
3.3 Evaluation of ObliQA
991 questions and their associated passages from the ObliQA dataset were evaluated by ADGM experts, as shown in Table 2. All questions were labeled as "Expected Question," confirming alignment with real-world compliance scenarios. The evaluation covered 1,847 passages, of which 1,605 (86.91%) were marked as relevant, 65 (3.52%) as irrelevant, and 8 (0.43%) as null. Additionally, 169 passages (9.14%) were identified as relevant but did not directly answer the corresponding questions. These findings highlight the high quality and relevance of the ObliQA dataset for research in regulatory compliance.
4 RePASs: Regulatory Passage Answer Stability Score
  • Every answer sentence must be supported by at least one sentence in the source passage(s).
  • The answer must cover all the obligations present in the source passages, meaning that all regulatory obligations should be reflected in the answer.
Step 1- Generation of Entailment and Contradiction NLI Pair Matrices: We generate an NLI pair matrix by comparing each passage sentence (premise) with each answer sentence (hypothesis) using the cross-encoder/nli-deberta-v3-xsmall He et al. (2021a). The model provides three outputs for each pair: probabilities for entailment, contradiction, and neutrality. These probabilities are organized into two matrices: the entailment matrix, which contains the entailment probabilities for each passage-answer sentence pair, and the contradiction matrix, which stores the contradiction probabilities for each sentence pair.
Step 2- Calculation of Entailment and Contradiction Scores: We reduce the size of the entailment and contradiction matrices by selecting the highest probability for each answer sentence across the passage sentences. Matrix reduction ensures that only the most relevant support from or conflict with the source passage is considered. The entailment score is computed by averaging the highest entailment probabilities across all answer sentences. Similarly, the contradiction score is obtained by averaging the highest contradiction probabilities.
$E_{s}=\frac{1}{N}\sum_{i=1}^{N}\max_{j}P_{\text{entailment}}(p_{j},a_{i})$ (1)
$C_{s}=\frac{1}{N}\sum_{i=1}^{N}\max_{j}P_{\text{contradiction}}(p_{j},a_{i})$ (2)
Step 3- Calculation of Obligation Coverage Score: The obligation coverage score evaluates how accurately the generated answer reflects the obligations present in the source passage(s). First, a dataset is created using an automated pipeline powered by the GPT-4-turbo-1106 model, which extracts and labels sentences from regulatory documents as either obligations or non-obligations using zero-shot learning. This labeled dataset is then used to fine-tune a LegalBERT-based classifier model Chalkidis et al. (2020), which is optimized for detecting obligatory sentences in regulatory texts. Both the source passages and generated answers are tokenized into sentences to enable a detailed, sentence-level comparison for obligation coverage.
The formula for the Obligation Coverage Score is given by:
Step 4- Calculation of the Regulatory Passage Answer Stability Score: The final step integrates the three components into a single composite score. The score penalizes contradictions and normalizes the result between 0 and 1.
$RePASs=\frac{\text{E}_{s}-\text{C}_{s}+\text{OC}_{s}+1}{3}$ (4)
4.1 Evaluation of RePASs
Metric Sub-metric Score
Standard Evaluation Metrics
SummaC - 0.5890
FactCC - 0.8571
BERTScore Precision 0.6710
Recall 0.7449
F1 0.7080
BLEU - 0.8300
ROUGE-1 Precision 0.8606
Recall 1.0000
F1 0.9218
ROUGE-2 Precision 0.8518
Recall 1.0000
F1 0.9155
ROUGE-L Precision 0.8606
Recall 1.0000
F1 0.9218
RePASs $E_{s}$ 0.7639
$C_{s}$ 0.3336
$OC_{s}$ 1.0000
RePASs 0.8101
Table 3: Performance of evaluation metrics on the gold standard dataset
Traditional metrics like BLEU Papineni et al. (2002), ROUGE Lin (2004), BERTScore Zhang et al. (2020), FactCC Kryscinski et al. (2020), and SummaC Laban et al. (2022) are used to evaluate text generation but fall short in RIRAG contexts. BLEU and ROUGE focus on n-gram overlap and lexical similarity, neglecting factual accuracy and the completeness of regulatory obligations. BERTScore captures semantic similarity but lacks tools to assess whether answers cover all obligations or avoid contradictions. While FactCC ensures factual consistency, it doesn’t verify that all necessary regulatory obligations are included. SummaC evaluates consistency with the source text but doesn’t account for regulatory adherence or contradictions. In contrast, RePASs is tailored specifically to regulatory compliance. It evaluates answers based on entailment, contradiction, and obligation coverage, ensuring that responses not only maintain coherence but also comply with regulatory requirements.
To assess RePASs, we tested it on real-world questions using expert-provided obligations and answers from ADGM. Table 3 shows that while ROUGE-L achieves a high F1 score (0.9218), it doesn’t ensure obligation coverage or the absence of contradictions. Similarly, BERTScore’s F1 score (0.7080) reflects semantic similarity but fails in regulatory compliance. FactCC’s score (0.8571) emphasizes factual consistency but doesn’t capture regulatory requirements. SummaC (0.5890) also falls short of the precision needed in these contexts. Table 3 illustrates that RePASs aligns well with expert judgments, and the results provide an indicative upper bound for the performance of an automated system.
5 Regulatory Information Retrieval and Answer Generation Task
The Regulatory Information Retrieval and Answer Generation (RIRAG) task includes two key subtasks: passage retrieval and answer generation. In the passage retrieval subtask, the system retrieves all relevant obligations from regulatory documents in response to a given question. The goal is to ensure that all pertinent obligations are identified and collected. In the answer generation subtask, the system synthesizes the retrieved information into a precise, informative response that comprehensively addresses the question. In this section, we describe a baseline approach for RIRAG, setting a standard that will support future research in the field.
5.1 Passage Retrieval
DRAGON+, SPLADEv2 and ColBERTv2 represent the three main architectures for neural retrieval, achieving SotA effectiveness with limited training resources and parameters (all around 110M parameters). NV-Embed-v2 and BGE-EN-ICL are the best-performing dense retrievers when the training resources are extensive and the parameters are largely scaled up (7.1B and 7.9B parameters, respectively). Additionally, BGE-EN-ICL supports In-Context Learning (ICL) Brown et al. (2020) for text embeddings, providing an option for obtaining task-adapted model with a low cost via adding example demonstrations in the model input.
Since the retrieval task in RIRAG searches forretrieve passages from a corpus of long documents, we also consider it as a Document-Aware Passage Retrieval (DAPR) Wang et al. (2024) task to jointly model the document context and passage retrieval. To this end, we apply rank fusion to linearly fuse the passage ranking by the neural or BM25 retrievers and the document ranking by the BM25 retriever. Wang et al. (2024) shows that this approach yields a consistent improvement on retrieval tasks across different domains. The details are available in Appendix A.7. The counterpart without rank fusion is denoted as passage-only.
For evaluating the retrieval module in RIRAG, we use recall@10 as the main evaluation metric, as we rely on the retrieval module to cover the relevant information as much as possible while leaving the burden of noise filtering to the answer-generation module. This approach is in line with LongRAG Jiang et al. (2024b). To additionally assess the ranking ability within the top-10 passages, MAP@10 (Mean Averaged Precision) is also used for diagnostic purposes.
For the experimental setup, we truncate the queries and the passages to 512 tokens. In rank fusion, we fuse the ranking of top-100 passages and the ranking of all the documents. The fusion weight on the document side is set to 0.1. The ranking scores are normalized (cf. Equation 5) with respect to the top-100 passages into the range of $[0,1]$ for filtering the passages conveniently in the answer-generation step. For BGE-EN-ICL, we consider the scenarios of zero-shot, 1-shot, 3-shot and 5-shot. The ICL examples are obtained via random sampling from the training set of ObliQA. According to their recommended usage, the instructions/query prefixes [^7] These models do not support instructions for the passage input. in NV-Embed-v2 and BGE-EN-ICLL are “Given a question, retrieve passages that answer the question” and “Given a financial-regulation query, retrieve relevant passages that answer the query”, respectively.
5.2 Answer Generation
The answer generation process begins once 10 relevant passages have been retrieved. In our post-retrieval stage, we implement a score-based filtering approach with a threshold of 0.2 to identify significant drops in relevance between consecutive passages. We also enforce a minimum score criterion of 0.7, ensuring that only the most relevant passages are considered for generating answers. Based on these filtered passages, we use the gpt-4-turbo-1106 model to generate comprehensive answers. The model follows a tailored prompt designed to simulate the role of a regulatory compliance assistant, integrating all relevant obligations and best practices from the passages into a cohesive response. The prompt reads:
5.3 Results
5.3.1 Passage Retrieval Results
The passage retrieval results on the test split of ObliQA are shown in Table 4. We find the models of DRAGON+, SPLADEv2 and ColBERTv2, which are fine-tuned with limited training data (i.e. MS MARCO only) and limited parameters, generalize poorly, achieving worse or on-par recall@10 and MAP@10 than simple BM25. ColBERTv2, the best model in this group, outperforms BM25 by 0.6 points MAP@10 and 1.1 points recall@10. On the other hand, the models of NV-Embed-v2 and BGE-EN-ICL, which are fine-tuned with more training data and more parameters, outperform BM25 significantly in all the settings. Compared with BM25, recall@10 and MAP@10 are improved by up to 3.4 points and up to 2.1 points, respectively. These results show that scaling the retrievers with more training data and parameters is necessary for good generalization on the retrieval task in RIRAG. We find ICL is important for adapting the neural models to this new task, while the advantage of including more examples in the demonstration is very marginal. Lastly, we find that rank fusion, which considers the document context during passage retrieval, improves the performance only marginally. We assume this is because the query generation process only considers the target passages as input, resulting in fairly self-contained query-evidence pairs.
Model Passage-only Rank fusion
R@10 M@10 R@10 M@10
BM25 76.1 62.4 76.4 62.5
DRAGON+ 74.3 58.1 74.5 58.2
SPLADEv2 75.7 60.4 75.9 60.7
ColBERTv2 77.7 63.0 77.7 63.2
NV-embed-v2 78.3 63.5 78.2 63.6
BGE-EN-ICL (0-shot) 77.1 62.4 77.4 62.5
BGE-EN-ICL (1-shot) 79.2 64.1 79.3 64.2
BGE-EN-ICL (3-shot) 79.5 64.1 79.5 64.2
BGE-EN-ICL (5-shot) 79.3 64.4 79.5 64.5
Table 4: Results of the retrieval task on the test dataset. R@10 and M@10 represent Recall@10 and MAP@10, respectively.
5.3.2 Answer Generation Results
Table 5 illustrates the performance differences between BM25(passage-only)+GPT-4 and BM25(rank fusion)+GPT-4 used in our answer generation task. Similar to the retrieval task results, where rank fusion showed slight performance improvements, we observe that BM25 (rank fusion) + GPT-4 outperforms BM25 (passage-only) + GPT-4 across all metrics. Specifically, rank fusion achieves higher Es (0.320 vs. 0.308), Cs (0.131 vs. 0.123), OCs(0.222 vs. 0.214), and RePASs (0.470 vs. 0.466). This highlights the connection between better passage retrieval and improved answer generation in regulatory tasks.
Method Es Cs OCs RePASs
BM25(passage-only)+GPT-4 0.308 0.123 0.214 0.466
BM25(rank fusion)+GPT-4 0.320 0.131 0.222 0.470
Table 5: Results of the answer generation task using RePASs on the ObliQA test dataset. Es, Cs, OCs and RePASs represent Entailment, Contradiction, Obligation Coverage and RePAS score, respectively.
6 Conclusion
In conclusion, this paper advances the field of RegNLP by introducing research infrastructure and metrics aimed at improving the overall precision of regulatory compliance. The automated question-passage generation framework and the ObliQA dataset contribute to the study of the retrieval and understanding of regulatory information. Furthermore, the regulatory information retrieval and answer generation task, supported by the novel RePASs evaluation metric, provides a framework for improving on our baseline approach.
7 Limitations
One significant limitation we encountered is the scarcity of publicly available datasets. Regulatory compliance often involves handling highly sensitive and confidential information, such as personal data or financial records, which presents challenges in terms of data privacy and security. To address this, we created a semi-synthetic dataset. During the validation phase, we faced difficulties stemming from multidisciplinary complexity and shortage of experts. However, experts evaluated a small subset of the dataset and confirmed that 100% of the generated questions met expectations, while 86% of the answer passages were deemed relevant to the questions.
Another key limitation is the lack of standardized formats for regulatory documents. Regulatory documents vary significantly across industries, regions, and even within the same organization, making it difficult to develop models capable of generalization.In this work, standardizing 40 documents required over 300 person-hours of effort. Additionally, the handling of tables and figures presented further challenges; tables were semi-automatically converted to text format, and figures were interpreted using the GPT-4 model.
Acknowledgement
Appendix A Appendix
A.1 Illustrative Example of RIRAG for Compliance Queries
Figure 2: An example of a comprehensive response by retrieving and synthesizing relevant passages from regulatory documents in response to a client question
Figure 2 illustrates a RIRAG task. In response to a client’s query about handling suspicious financial transactions, the system retrieves key regulations from several passages across the Anti-Money Laundering (AML) and General (GEN) sections of regulatory documents. By synthesizing these passages, the system generates a concise, comprehensive answer, streamlining the institution’s compliance with complex regulations.
A.2 Document Standardization and Structuring